# Anatomical Foundation Models for Brain MRIs

### v1

@@ Abstract
Deep Learning (DL) in neuroimaging has become increasingly relevant for detecting neurological conditions and neurodegenerative disorders. One of the most predominant biomarkers in neuroimaging is represented by brain age, which has been shown to be a good indicator for different conditions, such as Alzheimer’s Disease. Using brain age for pretraining DL models in transfer learning settings has also recently shown promising results, especially when dealing with data scarcity of different conditions. On the other hand, anatomical information of brain MRIs (e.g. cortical thickness) can provide important information for learning good representations that can be transferred to many downstream tasks. In this work, we propose AnatCL, an anatomical foundation model for brain MRIs that i.) leverages anatomical information with a weakly contrastive learning approach and ii.) achieves state-of-the-art performances in many different downstream tasks. To validate our approach we consider 12 different downstream tasks for diagnosis classification, and prediction of 10 different clinical assessment scores.
@@ 1 Introduction
Magnetic Resonance Imaging (MRI) is the foundation of neuroimaging, providing detailed views of brain structure and function that are crucial for diagnosing and understanding various neurological conditions. However, the complexity and high dimensionality of MRI data present significant challenges for automated analysis. Traditional machine learning methods often require extensive manual annotation, which is both time-consuming and expensive. Recently, contrastive deep learning techniques have emerged as powerful tools for extracting meaningful information from complex data, showing promise in improving the automated analysis of neuroimaging datasets.
Figure 1: We propose AnatCL, a contrastive approach that leverages anatomical information in brain MRIs in order to learn meaningful and general representation. AnatCL can transfer to different downstream tasks achieving comparable or better results that standard state-of-the-art approaches, which are based on brain age prediction.
@@
Contrastive learning methods, particularly those leveraging supervised contrastive loss functions, have been effective in creating models that learn robust representation spaces. These methods typically utilize metadata, such as patient age, to guide the learning process, with the aim of aligning similar data points (i.e. similar age) and distinguishing dissimilar ones in the representation space. While this approach has yielded promising results, it may fall short of fully capturing the multifaceted information inherent in MRI data and the anatomical complexity of the brain.
@@
For this reason, in this paper, we extend the capabilities of existing contrastive learning approaches by incorporating additional anatomical measures into the loss formulation. We propose AnatCL, a contrastive learning framework that builds upon previous works that leverages patient age as proxy metadata [12, 1] to also include information derived from anatomical measures (such as cortical thickness), computed on region of interests (ROIs) of common atlas such as the Desikan-Killiany parcellation [7] Specifically, we propose a modified contrastive learning loss function that integrates patient age with a targeted selection of three anatomical features: Mean Cortical Thickness, Gray Matter Volume, Surface Area. This selective integration was designed to assess the effectiveness of using potentially more informative anatomical features in conjunction with demographic data.
@@
We propose and test two versions of our framework: a local version that considers the distance between representations as the distance between local measurements in the Desikan format (a set of $68\times 3$ features), and a global version that considers the same three features across all Desikan regions ($3\times 68$ features). The local version aims to capture the localized anatomical variations within specific regions, while the global version aggregates these features across the entire brain to provide a more holistic view.
@@
An overview of AnatCL is provided by Fig. 1. First, we pretrain our models using the proposed method on the OpenBHB dataset, a large collection of healthy samples, then we test the resulting models on different downstream tasks across a variety of publicly available datasets (Sec. 3.1). Our results indicate that in some downstream tasks, our method can achieve comparable or better results than standard state-of-the-art approaches based on brain age prediction. These findings suggest that incorporating targeted anatomical features alongside patient age in contrastive learning frameworks can enhance the performance of neuroimaging models, leading to the development of models that learn more meaningful and generalizable representation spaces. Additionally, we also perform experiments to check whether anatomical information can be leveraged for predicting phenotypes and clinical assessments scores. To the best of our knowledge, this is one of the first works attempting this. In summary, our contribution is multifold:
@@
    We propose AnatCL, a foundation models based on both anatomical and demographic features to build robust and generalizeable representations of structural neuroimaging data;
@@
Furthermore, we also publicly release our pre-trained models for researchers worldwide at omittedlink.
@@ 2 Related Works
Machine learning on structural brain imaging.
@@
Current literature has been mainly devoted to segmentation tasks for brain anatomical data, such as brain tumor segmentation [2] or tissue segmentation [10], in a supervised setting requiring annotated maps. The deep neural network U-Net [25] and its extension nnU-Net [18] were very successful at such tasks. Another body of literature tackled phenotype and clinical prediction tasks from structural imaging, such as age prediction [14], Alzheimer’s disease detection [31], cognitive impairment detection, or psychiatric conditions prediction [22, 4]. Current approaches mostly involve traditional machine learning models such as SVM or penalized linear models trained in a supervised fashion to predict the targeted phenotypes. Since most studies only include a few hundred subjects in their datasets, larger models introducing more parameters (such as neural networks) do not necessarily translate into better performance [11, 26, 16], probably because of over-fitting.
Self-supervised learning on brain imaging.
@@
Self-supervised learning has been popularized by its impressive performance in Natural Language Processing (NLP) using auto-regressive models trained on web-scale text corpus and mainly deployed on generation tasks [8, 3]. Since it does not require a supervised signal for training, it is particularly appealing when large-scale data are available, and the annotation cost is prohibitive. Its applicability on neuroimage data is still unclear and very few studies have tackled this problem so far [17].  [12] demonstrated that contrastive learning integrating subjects meta-data can improve the classification of patients with psychiatric conditions.  [29] demonstrated that learning frameworks from NLP, such as sequence-to-sequence autoencoding, causal language modeling, and masked language modeling, can improve mental state decoding from brain activity recorded with functional MRI.
Contrastive Learning on brain imaging
@@
Recently, interests in contrastive learning approaches for brain age prediction has risen. Among different works we can identify y-Aware [12] and ExpW [1]. These works propose to utilize contrastive learning for continuous regression tasks, by employing a kernel to measure similarity between different samples. In these works, age is used as pre-training target, followed by transfer learning on specific diseases. However, they are limited to only considering age as pretraining feature. In this work, we extend this formulation to work with multiple attributes.
3 Materials and Methods
@@ 3.1 Experimental data
We use a collection of T1-weighted MRI scans comprising 7,908 different individuals, for a total of 21,155 images. The experimental data is gathered from different publicly available datasets, and we target 5 different neurological conditions: healthy samples (OpenBHB [13]), Alzheimer’s Disease (ADNI [23] and OASIS-3 [20]), schizophrenia (SchizConnect [30]) and Autism Spectrum Disorder (ABIDE I [9]). The overall composition of the cohorts used in this study is presented in Fig. 1B. Tab. 1 provides an overview of all the downstream tasks that we explore in this work.
OpenBHB
OpenBHB is a recently released dataset that aggregates healthy control (HC) samples from many public cohorts (ABIDE 1, ABIDE 2, CoRR, GSP, IXI, Localizer, MPI-Leipzig, NAR, NPC, RBP). Every scan comes from a different subject. We use OpenBHB as pretraining dataset for our method. Besides the structural scans and patients information, OpenBHB also provides 7 anatomical measures according to the Desikan-Kiliani parcellation [7] (cortical thickness mean and std, gray matter volume, surface area, integrated mean, gaussian curvature index and intrinsic curvature index [13]).
ADNI and OASIS-3
We include all phases ADNI-1, ADNI-2, ADNI-GO and ADNI-3 in our study, amounting to 633 HC, 712 MCI and 409 AD patients. For OASIS3 we include 685 patients respectively, containing 88 AD cases.
SchizConnect
We include anatomical MRIs for a total of 383 patients, divided in 180 HC, 102 with schizophrenia (broad), 74 with schizophrenia (strict), 11 schizoaffective patients and 9 with bipolar disorder.
ABIDE I
We include anatomical MRIs for a total of 1102 patients, divided in 556 HC, 339 patients with autism, 93 patients with Asperger’s Syndrome, and 7 with pervasive developmental disorder not otherwise specified (PDD-NOS).
Clinical assessment scores
In addition to predict the above diagnosis, we also consider some relevant clinical assessment included in the data, as reported in Tab. 1. For SchizConnect, we consider the Abormal Involuntary Movement Scale (AIMS) evaluated in three assessments (overall, uppder body, lower body), a Depression score based on the Calgary Scale, Handedness information, and GAIT measuremens with the Simpson-Angus-Scale (SAS). For ABIDE, we consider also handedness and three IQ scores, namely Full Scale IQ (FIQ), Visual IQ (VIQ) and Performance IQ (FIQ) measured with the Wechsler Abbreviated Scale (WASI).
@@
OpenBHB Age (HC)
Sex
ADNI Alzheimer’s Disease
@@
SchizConnect Schizophrenia Broad
Schizophrenia Strict
Bipolar Disorder
@@
ABIDE I Autism
Aspergers

### v3

@@ Abstract
Deep Learning (DL) in neuroimaging has become increasingly relevant for detecting neurological conditions and neurodegenerative disorders. One of the most predominant biomarkers in neuroimaging is represented by brain age, which has been shown to be a good indicator for different conditions, such as Alzheimer’s Disease. Using brain age for weakly supervised pre-training of DL models in transfer learning settings has also recently shown promising results, especially when dealing with data scarcity of different conditions. On the other hand, anatomical information of brain MRIs (e.g. cortical thickness) can provide important information for learning good representations that can be transferred to many downstream tasks. In this work, we propose AnatCL, an anatomical foundation model for brain MRIs that i.) leverages anatomical information in a weakly contrastive learning approach, and ii.) achieves state-of-the-art performances across many different downstream tasks. To validate our approach we consider 12 different downstream tasks for the diagnosis of different conditions such as Alzheimer’s Disease, autism spectrum disorder, and schizophrenia. Furthermore, we also target the prediction of 10 different clinical assessment scores using structural MRI data. Our findings show that incorporating anatomical information during pre-training leads to more robust and generalizable representations. Pre-trained models can be found at: https://github.com/EIDOSLAB/AnatCL.
@@ 1 Introduction
Magnetic Resonance Imaging (MRI) is the foundation of neuroimaging, providing detailed views of the brain structure and function that are crucial for diagnosing and understanding various neurological conditions. However, the complexity and high dimensionality of MRI data present significant challenges for automated analysis. Traditional machine learning methods often require extensive manual annotation, which is both time-consuming and expensive. Recently, contrastive deep learning techniques have emerged as powerful tools for extracting meaningful information from complex data with minimal requirement for annotations, showing promise in improving the automated analysis of neuroimaging datasets.
Figure 1: We propose AnatCL, a contrastive learning approach that leverages patient’s metadata (i.e. age) alongside anatomical information derived from structural MRIs, in order to learn meaningful and general representation. Given two input samples $x$ and $x_{i}$, the objective of AnatCL is to learn embeddings $z$ and $z_{i}$, such that their distance in the representation space is proportional to the difference in age and brain anatomy between the two samples.
@@
Contrastive learning methods, particularly those leveraging weakly supervised loss functions, have shown to be effective in creating models that learn robust representation spaces [12, 14]. These methods typically utilize metadata, such as the patient’s age, to guide the learning process, with the aim of aligning similar data points (i.e. similar age) and distinguishing dissimilar ones in the learned representation space. While this approach has yielded promising results, it may fall short of fully capturing the multifaceted information inherent in MRI data and the anatomical complexity of the brain, as it primarily relies on age alone.
@@
For this reason, in this paper, we extend the capabilities of weakly supervised contrastive learning approaches by incorporating additional anatomical measures into the loss formulation. We propose AnatCL, a contrastive learning framework that improves upon previous works leveraging patient age as proxy metadata [12, 1], to also include information derived from anatomical measures (such as cortical thickness), computed on region of interests (ROIs) of common atlas such as the Desikan-Killiany parcellation [7] Specifically, we propose a modified contrastive learning loss function that integrates patient age with a targeted selection of three anatomical features: Mean Cortical Thickness, Gray Matter Volume, Surface Area. This selective integration was designed to assess the effectiveness of using potentially more informative anatomical features in conjunction with demographic data.
@@
We propose and test two versions of our framework: a local version that considers the distance between local anatomical measurements in the Desikan format (a set of $68\times 3$ features), and a global version that considers the same three features across all Desikan regions ($3\times 68$ features). The local version aims to capture the localized anatomical variations between specific ROIs, while the global version aggregates these features across the entire brain to provide a more holistic view of the general anatomy. A graphical overview of AnatCL is provided in Fig. 1.
Figure 2: Overview of the datasets used in this work. We pre-train AnatCL on the large dataset OpenBHB of brain MRIs, and we test different downstream tasks on different conditions such as Alzheimer’s Disease, Schizophrenia, and Autism Spectrum Disorder (ASD).
@@
We pre-train our models using the proposed AnatCL loss on the OpenBHB dataset, a large collection of healthy samples, and then we test the resulting models on different downstream tasks across a variety of publicly available datasets (Sec. 4.1). Our results show that in several downstream tasks, our method can achieve better results than standard state-of-the-art approaches based on brain age prediction. These findings suggest that, in contrastive learning frameworks, incorporating anatomical features alongside the patient’s age can enhance the expressiveness of the learned representations, leading to the development of models that learn more meaningful and generalizable representation spaces. Additionally, we also perform experiments to assess whether anatomical information can be leveraged for predicting phenotypes and clinical assessment scores. To the best of our knowledge, this is one of the first works attempting to tackle this objective. In summary, our contribution is multifold:
@@
    We propose AnatCL, a foundation model based on both anatomical and demographic features to build robust and generalizable representations of structural neuroimaging data;
@@
Furthermore, we also publicly release our pre-trained models for researchers worldwide at https://github.com/EIDOSLAB/AnatCL.
@@ 2 Related Works
Many works have explored the potential of machine learning and deep learning in neuroimaging. Here, we review the most popular approaches, distinguishing relevant topics for our work. First, we start with standard machine learning approaches, then we explore self-supervised methods, and finally, we present state-of-the-art contrastive approaches for neuroimaging.
@@
Machine learning on structural brain imaging. Current literature has been mainly devoted to segmentation tasks for brain anatomical data, such as brain tumor segmentation [2] or tissue segmentation [10], in a supervised setting requiring annotated maps. The deep neural network U-Net [27] and its extension nnU-Net [19] were very successful at such tasks. Another body of literature tackled phenotype and clinical prediction tasks from structural imaging, such as age prediction [15], Alzheimer’s disease detection [33], cognitive impairment detection, or psychiatric conditions prediction [24, 4]. Current approaches mostly involve traditional machine learning models such as SVM or penalized linear models trained in a supervised fashion to predict the targeted phenotypes. Since most studies only include a few hundred subjects in their datasets, larger models introducing more parameters (such as neural networks) do not necessarily translate into better performance [14, 28, 17], probably because of over-fitting.
@@
Self-supervised learning on brain imaging. Self-supervised learning has been popularized by its impressive performance in Natural Language Processing (NLP) using auto-regressive models trained on web-scale text corpus and mainly deployed on generation tasks [8, 3]. Since it does not require a supervised signal for training, it is particularly appealing when large-scale data are available, and the annotation cost is prohibitive. Its applicability on neuroimage data is still unclear and very few studies have tackled this problem so far [18].  [12] demonstrated that contrastive learning integrating subjects meta-data can improve the classification of patients with psychiatric conditions.  [31] demonstrated that learning frameworks from NLP, such as sequence-to-sequence autoencoding, causal language modeling, and masked language modeling, can improve mental state decoding from brain activity recorded with functional MRI.
@@
Contrastive Learning on brain imaging Recently, interests in contrastive learning approaches for brain age prediction has risen. It has been demonstrated that contrastive learning performs better that traditional methods for neuroimaging [14]. Among different works we can identify y-Aware [12] and ExpW [1]. These works propose to utilize contrastive learning for continuous regression tasks, by employing a kernel to measure similarity between different samples. In these works, age is used as pre-training target, followed by transfer learning on specific diseases. However, they are limited to only considering age as pretraining feature. In this work, we extend this formulation to work with multiple attributes.
3 Method
In this section, we present the proposed method AnatCL. To introduce the adequate tools and concepts for understanding our work, we first provide an introduction to contrastive learning and weakly contrastive learning. Then, we present AnatCL.
3.1 Background
Figure 3: Graphical evaluation of AnatCL (local and global) on all tasks considered. Overall, AnatCL achieves higher results more consistently than any other baselines, with notable performance on ASD and Schizophrenia.
3.2 Weakly contrastive learning
To tackle the issue of continuous labels, weakly contrastive approaches employ the notion of a degree of “positiveness” between samples. This degree is usually defined by a kernel function $w_{i}=\mathcal{K}(y-y_{i})$, where $0\leq w_{i}\leq 1$ is computed by a Gaussian or a Radial Basis Function (RBF) kernel and $y,y_{i}$ are the continuous attributes of interest (e.g. age). The goal of weakly contrastive learning is thus to learn a parametric function $f:\mathcal{X}\rightarrow\mathbb{S}^{d}$ that maps samples with a high degree of positiveness ($w_{i}\sim 1$) close in the latent space and samples with a low degree ($w_{i}\sim 0$) far away from each other. y-Aware [12] proposes to do so by aligning positive samples with a strength proportional to the degree, i.e.:
$\mathcal{L}^{y-aware}=-\sum_{i}\frac{w_{i}}{\sum_{j}w_{j}}\log\left(\frac{\exp(sim(z,z_{i}))}{\sum_{t=1}^{N}\exp(sim(z,z_{t}))}\right)$ (1)
where $z=f(x)$, $z_{i}=f(x_{i})$, $z_{t}=f(x_{t})$, and $sim$ is a similarity function (e.g. cosine). A similar approach is also adopted by ExpW [1]. A limitation of this class of approaches is the reliance on one single attribute (age) to determine the alignment strength. They are thus not suited to leverage multiple attributes (e.g. anatomical measurements of the brain). Our proposed approach aims to solve this issue by extending the weakly contrastive formulation to include multiple attributes.
3.3 Anatomical Contrastive Learning
We propose a novel formulation for weakly contrastive learning to employ multiple anatomical attributes derived from the MRI scan, that we call AnatCL. The features we consider are derived from the standard Desikan-Kiliany parcellation [7], based on the information available in the OpenBHB dataset (see Sec. 4.1), composed of $K=68$ ROIs. Among the anatomical measurements, we consider cortical thickness (CT), gray matter volume (GMV), and surface area (SA) [13]. To include these measurements inside AnatCL, we propose two different formulation, global and local, based on how the similarity across two brain scans is computed.
3.3.1 Local Descriptor
Given a sample $x$, we consider the set of local descriptors $\Psi(x)=\{\psi(x)^{1},\dots\,\psi(x)^{K}\}$, where $\psi(x)^{k}\in\mathbb{R}^{3}$ for $k\in\{1,\dots,K\}$, and $K$ is the number of region of interest (ROIs) in the chosen atlas (Desikan). Each local descriptor is thus a vector composed of the three anatomical measurements CT, GMV, and SA, corresponding to a specific ROI of the atlas. In order to compute a degree of positiveness $\alpha_{i}$ based on the local descriptors of two samples $x$ and $x_{i}$, we consider the average cross-region similarity as:
$\alpha_{i}=\frac{1}{K}\sum_{k=1}^{K}s(\gamma(\psi^{k}),\gamma(\psi_{i}^{k}))$ (2)
where, for brevity $\psi^{k}=\psi(x)^{k}$ and $\psi_{i}^{k}=\psi(x_{i})^{k}$, $s$ can either be a similarity function or a kernel as in [12], and $\gamma(\psi)$ is a normalization function that normalizes each component of $\psi$ to a standard range $[0,1]$. In this work, for simplicity, we choose to employ cosine similarity as $s$. By plugging $a_{i}$ into Eq. 1, we obtain the local variant of AnatCL:
$\begin{split}\mathcal{L}_{AnatCL-L3}=\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\\ =-\sum_{i}\sum_{k=1}^{K}s(\gamma(\psi^{k}),\gamma(\psi_{i}^{k}))\log\left(\frac{\exp(sim(z,z_{i}))}{\sum_{t}^{N}\exp(sim(z,z_{t}))}\right)\end{split}$ (3)
where we omitted the normalization terms $1/\sum_{j}\alpha_{k}$ and $1/K$ for brevity.
3.3.2 Global Descriptor
Taking a complementary approach, we now consider global anatomy descriptors $\omega(x)^{j}\in\mathbb{R}^{K}$ for the entire brain with $j\in\{1,2,3\}$, that contain the values across all $K=68$ regions for each anatomical measurement. Similarly to above, in order to compute a degree of positiveness $\beta_{i}$ between two samples $x$ and $x_{i}$ given their global anatomical descriptors $\Omega(x)=\{\omega(x)^{1},\dots,\omega(x)^{3}\}$ and $\Omega(x_{i})=\{\omega(x_{i})^{1},\dots,\omega(x_{i})^{3}\}$ we compute the cross-measurement similarity (global descriptor) as:
$\beta_{i}=\frac{1}{3}\sum_{j=1}^{3}s(\omega^{j},\omega_{i}^{j})$ (4)
where, similarly to before, $\omega^{j}=\omega(x)^{j}$ and $\omega_{i}^{j}=\omega(x_{i})^{j}$. This time we do not need to normalize $\omega^{j}$ and $\omega_{i}^{j}$ as $s$ is evaluated between features of the same type, so the scale of the values is comparable. Also in this case, for the sake of simplicity, we employ a cosine similarity function for $s$. As before, by plugging $\beta_{i}$ in Eq. 1 we obtain the AnatCL global loss:
$\begin{split}\mathcal{L}_{AnatCL-G3}=\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\\ =-\sum_{i}\sum_{j=1}^{3}s(\omega^{j},\omega_{i}^{j})\log\left(\frac{\exp(sim(z,z_{i}))}{\sum_{t}^{N}\exp(sim(z,z_{t}))}\right)\end{split}$ (5)
where, as before, normalization terms are omitted.
3.3.3 Final objective function
The final objective function is a combination of our proposed AnatCL loss (either local or global) and the y-Aware loss [12] which considers age. Our final formulation is thus:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{AnatCL}+\lambda_{2}\mathcal{L}_{age}$ (6)
where $\mathcal{L}_{age}$ is defined as in Eq. 1 and $\mathcal{L}_{AnatCL}$ is defined by either Eq. 3 or Eq. 5. The positive weights $\lambda_{1}$ and $\lambda_{2}$ control the contribution of each loss term independently. It is worth noting that the considered anatomical features in AnatCL can be computed directly from the 3D MRI with standard tools such as FreeSurfer [^1] https://surfer.nmr.mgh.harvard.edu/, thus our method does not require any additional label in the dataset.
4 Experiments
In this section, we present the experimental data that we employ, and the experiments that we perform to assess the performance of AnatCL. We compare AnatCL to different baselines, composed of state-of-the-art approaches based on contrastive learning in several downstream tasks. An overview of downstream performance of AnatCL is provided in Fig. 3. From the plot, we can notice that AnatCL obtains state-of-the-art results in more tasks than any of the other methods.
@@ 4.1 Experimental data
We use a collection of T1-weighted MRI scans comprising 7,908 different individuals, for a total of 21,155 images. The experimental data is gathered from different publicly available datasets, and we target 5 different neurological conditions: healthy samples (OpenBHB [13]), Alzheimer’s Disease (ADNI [25] and OASIS-3 [22]), schizophrenia (SchizConnect [32]) and Autism Spectrum Disorder (ABIDE I [9]). The overall composition of the cohorts used in this study is presented in Fig. 2, while Tab. 1 provides an overview of all the downstream tasks that we explore in this work.
OpenBHB OpenBHB is a recently released dataset that aggregates healthy control (HC) samples from many public cohorts (ABIDE 1, ABIDE 2, CoRR, GSP, IXI, Localizer, MPI-Leipzig, NAR, NPC, RBP). Every scan comes from a different subject, for a total of 3984 MRI images. We use OpenBHB as the pre-training dataset for our method. Besides the structural scans and patients information, OpenBHB also provides 7 anatomical measures according to the Desikan-Kiliani parcellation [7] (cortical thickness mean and std, gray matter volume, surface area, integrated mean, gaussian curvature index and intrinsic curvature index [13]).
ADNI and OASIS-3 The Alzheimer’s Disease Neuroimaging Initiative (ADNI) is one of the largest and most popular datasets of MRIs for Alzheimer’s Disease (AD). In our study, we include all phases ADNI-1, ADNI-2, ADNI-GO and ADNI-3 amounting to 633 HC, 712 mild cognitive impairment (MCI) and 409 AD patients, for a total of 14,385 MRI images. We downloaded all raw acquisition in the DICOM format and performed VBM preprocessing with CAT12 [16]. We also collected all clinical information and processed it in order to distinguish between stable and progressive MCI. Additionally, we also employ the OASIS3 dataset, including 685 patients (88 AD cases and 597 control cases), for a total of 1354 additional images.
SchizConnect For schizophrenia we employ the SchizConnect, a large dataset of MRIs. We include structural MRIs for a total of 383 patients, divided into 180 HC, 102 with schizophrenia (broad), 74 with schizophrenia (strict), 11 schizoaffective patients, and 9 with bipolar disorder, for a total of 383 MRIs.
ABIDE I The Autism Brain Imaging Data Exchange (ABIDE) dataset is one of the most popular datasets of patients affected by conditions of Autism Spectrum Disorder (ASD). We leverage data from ABIDE-I, including structural MRIs for a total of 1102 patients, divided into 556 HC, 339 patients with autism, 93 patients with Asperger’s Syndrome, and 7 with pervasive developmental disorder not otherwise specified (PDD-NOS). The total number of images is 1102.
Clinical assessment scores and phenotypes In addition to predicting the above diagnosis, we also consider some relevant clinical assessments included in the data, as reported in Tab. 1. For SchizConnect, we consider the Abnormal Involuntary Movement Scale (AIMS) evaluated in three assessments (overall, upper body, lower body), a Depression score based on the Calgary Scale, Handedness information, and GAIT measurements with the Simpson-Angus-Scale (SAS). For ABIDE, we also consider handedness and three IQ scores, namely Full Scale IQ (FIQ), Visual IQ (VIQ), and Performance IQ (FIQ) measured with the Wechsler Abbreviated Scale (WASI). To the best of our knowledge, this is the first work attempting to tackle this objective.
@@
OpenBHB Age (HC), Sex
ADNI Alzheimer’s Diseas,
@@
SchizConnect Schizo. Broad & Strict,
Bipolar Disorder,
@@
ABIDE I Autism, Aspergers,